# Predictions of the Nancy Grace Roman Space Telescope Galactic Exoplanet Survey. IV. Lens Mass and Distance Measurements

### v1

@@ 1 Introduction
Requirement 1: Measure the mass function of exoplanets with $1M_{\oplus}<M<30M_{J}$ and $a\,{\geq}\,1$ AU to better than 15% per decade in mass.
Requirement 2: Measure the frequency of bound exoplanets with masses in the range $0.1M_{\oplus}<M<0.3M_{\oplus}$ to better than 25%.
Requirement 3: Determine the masses of, and distances to, host stars of 40% of the detected planets with a precision of < 20<\,20%.
Requirement 4: Measure the frequency of free floating planetary-mass objects from Mars to 10 Jupiter masses. If there is one $M_{\oplus}$ free-floating planet per star, measure this frequency to better than 25%.
Requirement 5: Estimate $\eta_{\oplus}$ (frequency of planets orbiting FGK stars with mass ratio and estimated projected semimajor axis within 20% of the Earth-Sun system) to a precision of 0.2 dex via extrapolation from larger and longer-period planets.
@@
Our current work is the fourth in a series of papers detailing the expected yields from RGES. Penny et al. (2019) estimated the yield of several Roman architectures for bound planets via microlensing. They used detailed simulations to assess the capabilities of a broad range of Roman mission architectures as well as a range of survey designs, and determined the expected microlensing detection efficiency and planet yield for each configuration. Ultimately, assuming a nominal Roman observing strategy of seven total fields observed in six seasons and a wide-passband cadence of ${\sim}15$ minutes, they predict a total yield of ${\sim}1400$ bound exoplanets with mass greater than ${\sim}0.1M_{\oplus}$.
@@
Figure 1: Left: GBTDS observational fields (black outlines) overlaid on the $K_{S}$-band extinction map from Surot et al. (2020). Right: On-sky locations for the selected sample of 3,000 simulated events, colored by $K_{S}$-band extinction. The color scale between the extinction map and simulated events is identical, therefore the single colobar can be shared.
@@ 2 Lens Mass & Distance Requirement
To place quantitative requirements on the GBTDS, one must make assumptions about the population of exoplanets being probed, including frequency as a function of mass and semimajor axis. Of course, this frequency is poorly known and is precisely what the survey is designed to constrain. Validating requirements therefore means determining whether if a given frequency is assumed the requirement will be met, regardless of the true frequency of planets and thus whether the requirement will in fact be met with the actual survey. At the time when the science requirements were first being formulated, the best estimate of the mass function of cold exoplanets detected by microlensing was that of Cassan et al. (2012). Penny et al. (2019) adopted a modified version of this mass function, assuming that the frequency of planets saturated a 2 planets per dex below a planet mass of $M_{\mathrm{p}}<5.2M_{\oplus}$,
@@
where $a$ is semi-major axis, $M_{\mathrm{p}}$ is planet mass, and $N$ is the average number of planets per star. Note that this is both a mass function and semimajor axis distribution, with the distribution of planets semimajor axis assumed to be uniform in $\log a$.
This form was used to define and verify science requirements EML 2.0.1, 2.0.2 Penny et al. (2019) and 2.0.4 Johnson et al. (2020), and will be used to verify EML 2.0.5. We note it was also used for the population of host planets for evaluating the yield of exomoons orbiting wide-separation planets from Roman by Lastovka et al. (2025). While there have since been many microlensing analyses which derive an updated MF or mass ratio function (e.g. Suzuki et al. (2016, 2018), Bennett et al. (2021), Zang et al. (2025)), it is essential that, when verifying the science requirements, we adopt the same assumptions as were used to define these requirements. We therefore adopt this modified version of the Cassan et al. (2012) MF when verifying EML 2.0.3 (Requirement 3) on the host star masses and distances (see Section 4.2).
@@ 2.1 Light Curve Higher-Order Effects
Typical binary light curve modeling parameters include $q$, the mass ratio between the two lens components and $s$, the separation between the two lens components, in units of the angular Einstein radius, $\theta_{E}$. The mass ratio and separation are almost always measured with reasonable precision, although there are well-known degeneracies that can make determining the true parameters difficult (e.g., Dominik (1999), Yee et al. (2021), Zhang et al. (2022)). The Einstein radius crossing time, $t_{E}$, is the time it takes for the lens-source relative motion to traverse the Einstein radius ($\theta_{E}$). This parameter is also typically well-measured, however in cases of faint sources with very high magnification planetary anomalies, it can be difficult to measure $t_{E}$ precisely because this parameter must be measured from the low-magnification region of light curves (Alard, 1997). Another typical light curve modeling parameter is the de-reddened source star magnitude and color ($I_{S,0},(V_{S,0}-I_{S,0})$). This can be used to estimate the source star angular radius, $\theta_{*}$ (Kervella et al., 2004; Boyajian et al., 2014; Adams et al., 2018).
@@
While some of these routine observables, namely the dimensionless quantities $q$ and $s$, provide insight into the properties of the detected planetary systems, it is generally quite valuable if physical quantities can also be measured, such as the planet mass $M_{\mathrm{p}}$ (in, e.g,. $M_{\oplus}$ or $M_{J}$, the host star mass $M_{L}$ (in $M_{\odot}$), the semimajor axis of the planet $a$ (in au), as well as the distance to the system $D_{L}$ (in pc or kpc). Importantly, when determined as a function of these parameters, microlensing demographic constraints can then be directly compared to those determined with other methods. Generally, translating the routine light curve observables to physical parameters requires a measurement of the host mass $M_{L}$ and distance $D_{L}$. We note that, with a measurement of these two parameters, it is possible to express $s$ in physical units, i.e., the instantaneous projected separation $a_{\perp}=s\theta_{\rm E}D_{L}$. However, the true semimajor axis is related to $a_{\perp}$ through the phase, inclination, eccentricity, and argument of periastron of the orbit, which are generally unknown.
@@
In order to achieve direct mass measurements of the host and companion using only information from the microlensing light curve itself, one must precisely measure two quantities from higher-order effects in the microlensing light curve. Finite source effects occur when the light curve deviates from a point-source magnification due to the finite size of the source in regions where the magnification changes quickly over the size of the source. This effect is prominent for sources that approach within a few source radii from a caustic curve. The magnitude of finite source effects depends on the angular size of the source ($\theta_{*}$), or sometimes expressed in units of the angular Einstein ring radius ($\rho$), where $\rho=\theta_{*}/\theta_{E}$. Further, the source radius crossing time, $t_{*}$ is another measure of the finite source effect. Many planetary microlensing events to date have had $t_{*}$ measured reasonably well, which allows another way to measure the angular Einstein radius; $\theta_{E}=t_{E}\,\theta_{*}/t_{*}$. Notice a measurement of the finite source effect allows one to infer $\theta_{E}$, since $\theta_{*}$ can typically be inferred from the source flux and color with a measurement or assumption of the extinction. The microlensing parallax, $\pi_{E}$, can be measured by a distortion in the light curve due to the acceleration of the observer, typically due to the orbital motion of the Earth (Alcock et al., 1995). Microlensing parallax is always present in a microlensing event, but not always measurable. If measured, usually the East component is more tightly constrained than the North component because the orbital acceleration of the Earth perpendicular to the line-of-sight to the galactic bulge is largely in the East-West direction.
@@
$M_{L}=\frac{c^{2}}{4G}\theta_{E}^{2}\frac{D_{S}D_{L}}{D_{S}-D_{L}},$ (2)
@@
$M_{L}=\frac{c^{2}}{4G}\frac{\textrm{AU}}{{\pi_{E}}^{2}}\frac{D_{S}-D_{L}}{D_{S}D_{L}}$ (3)
Therefore, equations 2 and 3 can be combined to yield the lens mass, with no dependence on the lens or source distances:
$M_{L}=\frac{c^{2}}{4G}\theta_{E}\frac{\textrm{AU}}{\pi_{E}}$ (4)
So, when both $\theta_{E}$ and $\pi_{E}$ are well-measured it is possible to calculate the complete solution to a microlensing event. This is largely driven by the fact that unambiguous measurements of $\pi_{E}$ and $\theta_{E}$ lead to a geometric measurement of $M_{L}$ and $D_{L}$, which avoids dependencies on somewhat complex astrophysics (metallicity-dependent, extinction, stellar atmosphere, etc). One exception is the inference of $\theta_{*}$ from the flux and color of the source, which depends on knowing the extinction and the color-surface brightness relation. This relation requires accurate and empirical-based calibrations using known stars, typically of solar-type (Boyajian et al., 2014). See also Equation 11 which shows the relation for lens distance ($D_{L}$).
@@ 2.2 Measuring Lens Flux with Roman
In addition to the higher-order effects that can be measured from the light curve model fitting, there is a third quantity that can yield a mass-distance relation. A measurement of the lens star flux can be used alongside empirical mass-luminosity relations (Henry & McCarthy, 1993; Kenyon & Hartmann, 1995; Henry et al., 1999; Delfosse et al., 2000) or theoretical isochrones (Girardi et al., 2002; Bressan et al., 2012) to yield a third lens mass-distance relation:
${\cal F}_{\lambda}=\frac{C_{\lambda}f_{\lambda}(M_{L})}{D_{L}^{2}}\ ,$ (5)
where the relationship between the lens flux (${\cal F}_{\lambda}$) and distance ($D_{L}$) depends on the mass-luminosity relation, $f_{\lambda}(M_{L})$, and the calibration parameter, $C_{\lambda}$, both of which are passband-dependent. Isolating the flux from the lens typically requires high-resolution imaging of the target(s) to disambiguate any nearby unrelated and/or blended stars from the true microlensing target. An important fitting parameter in the microlensing modeling is the source flux ($f_{S}$) in units of flux or magnitude in a given filter. One can then compare the measurement of the target brightness in the high-resolution imaging to that of the source brightness from the light curve model. The difference in these two brightnesses is often referred to as “excess flux” at the position of the source, which can be attributed to the lens star, a companion to the source or lens, or an unrelated ambient field star that happens to be blended with the target(s). Excess flux measurements have been made for several microlensing targets in the past; Bennett et al. (2010); Kubas et al. (2012); Koshimoto et al. (2017a).
@@
$\vec{\mu}_{\textrm{rel,H}}=\vec{\mu}_{\textrm{rel,G}}+\frac{{\nu_{\earth}}{\pi_{\textrm{rel}}}}{AU}\ ,$ (8)
where $\nu_{\earth}$ is Earth’s projected velocity relative to the Sun at the time of peak magnification in the Geocentric reference frame and $\pi_{\rm{rel}}\equiv\rm{AU}/D_{L}-\rm{AU}/D_{S}$. For distant lenses (> 4>\,4 kpc) toward the galactic bulge, the difference in relative proper motion between these reference frames is negligible.
@@ 2.3 Measuring Lens-Source Proper Motion with Roman
We also estimate the fractional error on the relative proper motion, $\sigma_{\mu_{\rm rel}}/\mu_{\rm rel}$, using the CDCS method, which measures the displacement of the centroid between different filters, arising from the distinct spectral energy distribution of the source and lens (e.g. different color stars). For each filter combination, we compute the error in proper motion by propogating the error in measuring the centroid in the two filters which is given by:
@@ 2.4 Euclid and HST Precursor Imaging of the GBTDS Fields
The prospect of Euclid to deliver microlensing demographics has been considered in several past studies (Beaulieu et al., 2010; Penny et al., 2013). With the launch of Euclid, additional work has been conducted that includes studying how Roman microlensing mass measurements can be improved with early Euclid observations (Bachelet et al., 2022). Further, Bachelet & Penny (2019) and Beaulieu & Bachelet (2021) investigated science cases that can be optimized with simultaneous Euclid and Roman observations later this decade. In March 2025, Euclid obtained high-resolution images covering 5.2 deg2 of the inner bulge south of the GC (including the GBTDS footprint) with the VIS Camera. Compared to the regular Euclid observing sequence, the exposure time was slightly shorter (400 sec instead of 550 sec), with 16 dithers at each of the nine total pointings to ensure a well-sampled PSF. This program is known as the Euclid Galactic Bulge Survey (EGBS). Additional calibration fields were observed under the same thermal environment before and after the EGBS, which ensures a proper PSF modeling. The data is currently undergoing processing by the Euclid Exoplanet Science Working Group, with a scheduled public release in June 2026.

### v2

@@ 1 Introduction
Requirement 1 (EML 2.0.1): Measure the mass function of exoplanets with $1M_{\oplus}<M<30M_{J}$ and $a\,{\geq}\,1$ AU to better than 15% per decade in mass.
Requirement 2 (EML 2.0.2): Measure the frequency of bound exoplanets with masses in the range $0.1M_{\oplus}<M<0.3M_{\oplus}$ to better than 25%.
Requirement 3 (EML 2.0.3): Determine the masses of, and distances to, host stars of 40% of the detected planets with a precision of < 20<\,20%.
Requirement 4 (EML 2.0.4): Measure the frequency of free floating planetary-mass objects from Mars to 10 Jupiter masses. If there is one $M_{\oplus}$ free-floating planet per star, measure this frequency to better than 25%.
Requirement 5 (EML 2.0.5): Estimate $\eta_{\oplus}$ (frequency of planets orbiting FGK stars with mass ratio and estimated projected semi-major axis within 20% of the Earth-Sun system) to a precision of 0.2 dex via extrapolation from larger and longer-period planets.
@@
Our current work is the fourth in a series of papers detailing the expected yields from RGES. Penny et al. (2019) estimated the yield of several Roman architectures for bound planets via microlensing. They used detailed simulations to assess the capabilities of a broad range of Roman mission architectures as well as a range of survey designs, and determined the expected microlensing detection efficiency and planet yield for each configuration. Ultimately, assuming a nominal Roman observing strategy of seven total fields observed in six seasons, a wide-passband cadence of ${\sim}15$ minutes, and a fiducial mass function based on the results from Cassan et al. (2012), they predict a total yield of ${\sim}1400$ bound exoplanets with mass greater than ${\sim}0.1M_{\oplus}$.
@@
Figure 1: Left: GBTDS observational fields (black outlines) overlaid on the $K_{S}$-band extinction map from Surot et al. (2020). Right: On-sky locations for the selected sample of 3,000 simulated events, colored by $K_{S}$-band extinction. The color scale between the extinction map and simulated events is identical, therefore the single colorbar can be shared.
@@ 2 Lens Mass & Distance Requirement
To place quantitative requirements on the GBTDS, one must make assumptions about the population of exoplanets being probed, including their frequency as a function of mass and semi-major axis. Of course, this frequency is poorly known and is precisely what the survey is designed to constrain. Validating requirements therefore means determining whether if a given frequency is assumed the requirement will be met, regardless of the true frequency of planets and thus whether the requirement will in fact be met with the actual survey. At the time when the science requirements were first being formulated, the best estimate of the mass function of cold exoplanets detected by microlensing was that of Cassan et al. (2012). Penny et al. (2019) adopted a modified version of this mass function, assuming that the frequency of planets saturated at two planets per dex below a planet mass of $M_{\mathrm{p}}<5.2M_{\oplus}$,
@@
where $a$ is semi-major axis, $M_{\mathrm{p}}$ is planet mass, and $N$ is the average number of planets per star. Note that this is both a mass function and semi-major axis distribution, with the distribution of planets semi-major axis assumed to be uniform in $\log a$.
This form was used to define and verify science requirements EML 2.0.1, 2.0.2 (Penny et al., 2019) and 2.0.4 (Johnson et al., 2020), and will be used to verify EML 2.0.5. We note it was also used for the population of host planets for evaluating the yield of exomoons orbiting wide-separation planets from Roman by Lastovka et al. (2025). While there have since been many microlensing analyses which derive an updated MF or mass ratio function (e.g. Suzuki et al. (2016, 2018), Bennett et al. (2021), Zang et al. (2025)), it is essential that, when verifying the science requirements, we adopt the same assumptions as were used to define these requirements. We therefore adopt this modified version of the Cassan et al. (2012) MF when verifying EML 2.0.3 (Requirement 3) on the host star masses and distances (see Section 4.2).
@@ 2.1 Light Curve Higher-Order Effects
Following the formalism of Gould (2000a), we can define the angular Einstein radius ($\theta_{E}$) as the angular radius of the Einstein ring as seen by an observer. It can be expressed as follows:
$\theta_{E}\equiv\sqrt{\kappa M_{L}\pi_{\rm rel}}$ (2)
where $M_{L}$ is the mass of the lens, $\kappa=4G/c^{2}$AU, and $\pi_{\rm rel}\equiv\pi_{L}-\pi_{S}$ is the relative lens-source parallax. It is also useful to define the microlensing parallax vector $\vec{\pi}_{E}$, which has a magnitude
$\pi_{E}=\frac{\pi_{\rm rel}}{\theta_{E}},$ (3)
and direction equal to that of the relative lens-source proper motion $\vec{\mu}_{\rm rel}$. Lastly, the relative lens-source proper motion has a magnitude equal to $\mu_{\rm rel}=\theta_{E}/t_{E}$ (see also Eq. 11).
@@
While the above formulae are defined for a single lens and single source, typical binary light curve modeling includes additional parameters $q$, the mass ratio between the two lens components and $s$, the separation between the two lens components, in units of the angular Einstein radius, $\theta_{E}$. The mass ratio and separation are almost always measured with reasonable precision, although there are well-known degeneracies that can make determining the true parameters difficult (e.g., Dominik (1999), Yee et al. (2021), Zhang et al. (2022)). The Einstein radius crossing time, $t_{E}$, is the time it takes for the lens-source relative motion to traverse the Einstein radius ($\theta_{E}$). This parameter is also typically well-measured, however in cases of faint sources with very high magnification planetary anomalies, it can be difficult to measure $t_{E}$ precisely because this parameter must be measured from the low-magnification region of light curves (Alard, 1997). Another typical light curve modeling parameter is the de-reddened source star magnitude and color (e.g. $I_{S,0},(V_{S,0}-I_{S,0})$). This can be used to estimate the source star angular radius, $\theta_{*}$, using the dereddened flux of the source and color-surface brightness relations (Kervella et al., 2004; Boyajian et al., 2014; Adams et al., 2018).
@@
While some of these routine observables, namely the dimensionless quantities $q$ and $s$, provide insight into the properties of the detected planetary systems, it is generally quite valuable if physical quantities can also be measured, such as the planet mass $M_{\mathrm{p}}$ (in, e.g,. $M_{\oplus}$ or $M_{J}$), the host star mass $M_{L}$ (in $M_{\odot}$), the semi-major axis of the planet $a$ (in au), as well as the distance to the system $D_{L}$ (in pc or kpc). Importantly, when determined as a function of these parameters, microlensing demographic constraints can then be directly compared to those determined with other methods. Generally, translating the routine light curve observables to physical parameters requires a measurement of the host mass $M_{L}$ and distance $D_{L}$. We note that, with a measurement of these two parameters, it is possible to express $s$ in physical units, i.e., the instantaneous projected separation $a_{\perp}=s\theta_{\rm E}D_{L}$. However, the true semi-major axis is related to $a_{\perp}$ through the phase, inclination, eccentricity, and argument of periastron of the orbit, which are generally unknown.
@@
In order to achieve direct mass measurements of the host and companion using only information from the microlensing light curve itself, one must precisely measure two quantities from higher-order effects in the microlensing light curve. Finite source effects occur when the light curve deviates from a point-source magnification due to the finite size of the source in regions where the magnification changes quickly over the size of the source. This effect is prominent for sources that approach within a few source radii from a caustic curve. The magnitude of finite source effects depends on the angular size of the source ($\theta_{*}$), in units of the angular Einstein ring radius ($\rho$), where $\rho=\theta_{*}/\theta_{E}$. Further, the source radius crossing time, $t_{*}\equiv\rho\,t_{E}$, is another measure of the finite source effect. Many planetary microlensing events to date have had $t_{*}$ measured reasonably well, which allows another way to measure the angular Einstein radius; $\theta_{E}=t_{E}\,\theta_{*}/t_{*}$. Notice a measurement of the finite source effect allows one to infer $\theta_{E}$, since $\theta_{*}$ can typically be inferred from the source flux and color with a measurement or assumption of the extinction. As mentioned previously the microlensing parallax, $\pi_{\rm E}$, can be measured by a distortion in the light curve due to the acceleration of the observer, typically due to the orbital motion of the Earth (Alcock et al., 1995). Microlensing parallax is always present in a microlensing event, but not always measurable. If measured, usually the East component is more tightly constrained than the North component because the orbital acceleration of the Earth perpendicular to the line-of-sight to the galactic bulge is largely in the East-West direction.
@@
$M_{L}=\frac{\theta_{E}^{2}}{\kappa\pi_{\rm rel}}=\frac{c^{2}}{4G}\theta_{E}^{2}\frac{D_{S}D_{L}}{D_{S}-D_{L}},$ (4)
@@
$M_{L}=\frac{\pi_{\rm rel}}{\kappa\pi_{E}^{2}}=\frac{c^{2}}{4G}\frac{\textrm{AU}^{2}}{{\pi_{E}}^{2}}\frac{D_{S}-D_{L}}{D_{S}D_{L}}$ (5)
Therefore, equations 4 and 5 can be combined to yield the lens mass, with no dependence on the lens or source distances:
$M_{L}=\frac{\pi_{E}}{\kappa\theta_{E}}=\frac{c^{2}\textrm{AU}}{4G}\frac{\theta_{E}}{\pi_{E}}.$ (6)
Similarly, the relative lens-source parallax is given by
$\pi_{\rm rel}=\theta_{E}\pi_{E},$ (7)
and the lens distance is given by
$D_{L}^{-1}=\frac{\theta_{E}\pi_{E}}{\textrm{AU}}+D_{S}^{-1}.$ (8)
Thus, when both $\theta_{E}$ and $\pi_{E}$ are well-measured, it is possible to measure the mass of the lens, and with an estimate or measurement of the distance to the source, calculate the lens distance as well and thus obtain complete solution to a properties of the microlensing lens system in physical units. This is largely driven by the fact that unambiguous measurements of $\pi_{E}$ and $\theta_{E}$ lead to a geometric measurement of $M_{L}$ and $D_{L}$, which avoids dependencies on somewhat complex astrophysics (metallicity-dependent, extinction, stellar atmosphere, etc). One exception is the inference of $\theta_{*}$ from the flux and color of the source, which depends on knowing the extinction and the color-surface brightness relation. This relation requires accurate and empirical-based calibrations using known stars, typically of solar-type (Boyajian et al., 2014).
@@ 2.2 Measuring Lens Flux with Roman
In addition to the higher-order effects that can be measured from the light curve model fitting, there is a third quantity that can yield a mass-distance relation. A measurement of the lens star flux in a given passband ${\cal F}_{\lambda}$ can be used alongside a mass-luminosity relationship in that passband (either from empirical mass-luminosity relations (Henry & McCarthy, 1993; Kenyon & Hartmann, 1995; Henry et al., 1999; Delfosse et al., 2000) or theoretical isochrones (Girardi et al., 2002; Bressan et al., 2012)) to yield a third lens mass-distance relation
${\cal F}_{\lambda}=\frac{L_{\lambda}(M_{L})}{4\pi D_{L}^{2}}\$(9)
Here $L_{\lambda}(M_{L})$ is the relationship between the mass of a star and its specific luminosity, i.e., its luminosity in a given filter or passband. It is worthwhile noting that this relationship depends not only on the specific passband, but in principle also depends on other properties of the star such as its metallicity, surface gravity, age, etc., which are generally unknown. Isolating the flux from the lens typically requires high-resolution imaging of the target(s) to disambiguate any nearby unrelated and/or blended stars from the true microlensing target. An important fitting parameter in the microlensing modeling is the source flux ($f_{S}$) in units of flux or magnitude in a given filter. One can then compare the measurement of the target brightness in the high-resolution imaging to that of the source brightness from the light curve model. The difference in these two brightnesses is often referred to as “excess flux” at the position of the source, which can be attributed to the lens star, a companion to the source or lens, or an unrelated ambient field star that happens to be blended with the target(s). Excess flux measurements have been made for several microlensing targets in the past; Bennett et al. (2010); Kubas et al. (2012); Koshimoto et al. (2017a).
@@
$\vec{\mu}_{\textrm{rel,H}}=\vec{\mu}_{\textrm{rel,G}}+\frac{{\vec{v}_{\earth}}{\pi_{\textrm{rel}}}}{AU}\ ,$ (12)
where $\vec{v}_{\earth}$ is Earth’s projected velocity relative to the Sun at the time of peak magnification in the Geocentric reference frame and $\pi_{\rm{rel}}\equiv\rm{AU}/D_{L}-\rm{AU}/D_{S}$. For distant lenses (> 4>\,4 kpc) toward the galactic bulge, the difference in relative proper motion between these reference frames is negligible.
@@ 2.3 Measuring Lens-Source Proper Motion with Roman
We also estimate the fractional error on the relative proper motion, $\sigma_{\mu_{\rm rel}}/\mu_{\rm rel}$, using the CDCS method, which measures the displacement of the centroid between different filters, arising from the distinct spectral energy distribution of the source and lens (e.g. different color stars). For each filter combination, we compute the error in proper motion by propagating the error in measuring the centroid in the two filters which is given by:
@@ 2.4 Euclid and HST Precursor Imaging of the GBTDS Fields
The prospect of Euclid to deliver microlensing demographics has been considered in several past studies (Beaulieu et al., 2010; Penny et al., 2013). With the launch of Euclid, additional work has been conducted that includes studying how Roman microlensing mass measurements can be improved with early Euclid observations (Bachelet et al., 2022). Further, Bachelet & Penny (2019) and Beaulieu & Bachelet (2021) investigated science cases that can be optimized with simultaneous Euclid and Roman observations later this decade. In March 2025, Euclid obtained high-resolution images covering 5.2 deg2 of the inner bulge south of the GC (including the GBTDS footprint) with the VIS Camera. Compared to the regular Euclid observing sequence, the exposure time was slightly shorter (400 sec instead of 550 sec), with 16 dithers at each of the nine total pointings to ensure a well-sampled PSF. This program is known as the Euclid Galactic Bulge Survey (s). Additional calibration fields were observed under the same thermal environment before and after the EGBS, which ensures a proper PSF modeling. The data is currently undergoing processing by the Euclid Exoplanet Science Working Group, with a scheduled public release in June 2026.